Operations Report – November 28th

Crew 216 Operations Report 28-Nov-19
SOL: 4
Name of person filing report: Mike Lawson
Non-nominal systems: RAM rover power charging stations inoperable, electrical issue, ongoing.
Notes on non-nominal systems:
Generator: run
Hours run: 12.3
From what time last night: 1902
To what time this morning: 0723
List any additional daytime hours when the generator was run: N/A Solar— SOC % (Before generator is run at night): 69 %
Diesel Reading – 40
Station Propane Reading – 58
Ethanol Free Gasoline: N/A
Water (loft tank) (gal): 47
Water Meter (units): 1462347
Water (static tank) (gal): 375
Static to Loft Pump used – yes
Water in Green Hab (gal): 267
Water in Science Dome (gal): 0
Toilet tank emptied: yes
Deimos rover used: no
Beginning charge:
Ending charge:
Currently charging:
Sojourner rover used:
Beginning charge:
Ending charge:
Currently charging:
Spirit rover used: no
Hours: 114.5
Beginning charge: (Before EVA): 100
Ending charge: (On return from EVA, before recharging): 100
Currently charging:
Opportunity rover used: no
Hours: 68.7
Beginning charge: 100
Ending charge: 100
Currently charging: yes
Curiosity rover used: no
Hours: 119.6
Beginning charge: 100
Ending charge: (On return from EVA, before charging): 100
Currently charging: yes
Notes on rovers: will have to juggle charging sequence.
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
# Hours the ATVs were used today: 0
Notes on ATVs:
HabCar used and why, where?
CrewCar used and why, where?
General notes and comments:
Summary of internet: No serious issues.
Summary of suits and radios: Power strip to radio charger farm was inadvertently switched off, setting low power alarm off on one radio. Reset, no further issues.
Summary of Hab operations: Routine maintenance and housekeeping. Summary of GreenHab operations: Routine watering.
Summary of Science Dome operations: None, crew holiday.
Summary of RAM operations: RAM Rover power charging stations inop, electrical issues remain unresolved. RAM used to fabricate two baking pans from residue food cans for use in preparing holiday meal. Rover maintenance EVA performed.
Summary of any observatory issues: None, crew holiday.
Summary of health and safety issues: None.
Questions, concerns and requests to Mission Support: 1 bag NB trash in rear airlock. Request that a knife sharpener be added to the kitchen set for future mission. Crew used a field rock for sharpening. expedient (obsidian sample) for immediate need.

Copyright © The Mars Society. All rights reserved. | Main Site